OpenStudy (anonymous):

One million atoms of a particular element have a mass of 1.99x10ˆ-17 grams. How many moles of this element are there? What is the molar mass of this element? What is the identity of this element?

OpenStudy (chmvijay):

1,000, 000 atoms has mass of 1.99*10^-17 6.023 *10^23 atoms has mass of = 1.99*10^-17 * 1000000 / 6.023*10^23 cn u solve the above maths avogardro no makes one mole of that element and it is the atomic mass of that element

OpenStudy (anonymous):

Thanks chmvijay! Let me start on this one and I let you know if I have any other questions!

OpenStudy (anonymous):

Ok, I'm getting a negative number. What am I doing wrong?

OpenStudy (joannablackwelder):

Make sure the -17 is an exponent. Can you write what you have done?

OpenStudy (anonymous):

(1.99x10^-17) x 1000000 = 1.99x10^-11 (1.99x10^-11) / (6.023x10^23) = 3.304001328x10^-35